40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Aug 22 18:58:12 2012, Jenne, Update, PEM, Weird BLRMS increase 10x
    Reply  Thu Aug 23 00:18:54 2012, Jenne, Update, PEM, Weird BLRMS increase BS_OSEMsensors_higherFreqPeakIsOlder_LowerFreqPeakIsMoreRecent.pdf
    Reply  Thu Aug 23 10:08:13 2012, Steve, Update, PEM, seismometers? bajaMag3.pngseisvssus.png
Message ID: 7253     Entry time: Thu Aug 23 00:18:54 2012     In reply to: 7251
Author: Jenne 
Type: Update 
Category: PEM 
Subject: Weird BLRMS increase 
While I was gone for dinner break, the BLRMS went back to normal. Then, almost 2 hours later, another peak appeared, this time closer to 1Hz. Den noticed that it was hard to maintain any lock, since the optics were ringing up so much.

The MC was moving pretty significantly, and just to check, I turned off the WFS for a moment. The MC transmitted power was fluctuating by almost 50% until I turned the WFS back on.

Attached is a spectrum of the BS OSEM sensors. The higher frequency peak around 1.65Hz is from the time I posted the time series about earlier. The lower frequency peak around 1.15Hz is from the second interval of noise.

Now, the noise is gone, and things are back to normal (for now....)
Attachment 1: BS_OSEMsensors_higherFreqPeakIsOlder_LowerFreqPeakIsMoreRecent.pdf  98 kB  | Show | Show all
ELOG V3.1.3-